EXPERIMENTAL THERMOCHEMISTRY OF GAS PHASE CYTOSINE TAUTOMERS

\underline{A. M. MORRISON} and G. E. DOUBERLY, DEPARTMENT OF CHEMISTRY, UNIVERSITY OF GEORGIA, ATHENS, GEORGIA 30602-2556.

Enthalpies of interconversion are measured for the three lowest energy tautomers of isolated cytosine. The equilibrium distribution of tautomers near 600 K is frozen upon the capture of the gas phase species by low temperature helium nanodroplets. The temperature dependence of the gas phase cytosine tautomer populations is determined with infrared laser spectroscopy of the helium solvated species. The interconverison enthalpies obtained from the van’t Hoff relation are 1.14 ± 0.21 and 1.63 ± 0.12 for the C31 ⇀ ↽ C32 and C31 ⇀ ↽ C1 equilibria, respectively. C31 and C32 are rotamers of an enol tautomer, and C1 is a keto tautomer. The interconversion enthalpies are compared to recent CCSD(T) thermochemistry calculations of cytosine tautomers.